OpenStudy (anonymous):

To date, Jay Ward has a cumulative earnings of $95,000. This week he is paid $3,000. The total amount of Social Security Tax for this week is: (Assume a rate of 6.2% on $97,500 for Social Security and 1.45% for Medicare).

OpenStudy (anonymous):

\[\text{NetPay}=3000-(((95000+3000)-97500)0.062)-(3000*.0145)=\text{$\$$2925}.50\]

OpenStudy (anonymous):

The question I have is do you include what the employer would pay as well as what the employee would pay? If not, the Social Security Tax would be $155. If so, it would be double, or $310. $97,500 Social Security base -95,000 Cumulative Earnings =2,500 $2,500 * 6.2% (0.62) = $155

OpenStudy (anonymous):

My response was from the view of the employee when she opened her payroll check. The problem statement does not refer to the employer's payroll taxes. Your calculation on the Social Security Tax is correct and mine was incorrect. Jay's gross pay did exceed the maximum taxable SS gross by $500, that is, only $2500 was subject to the SS deduction. Do not award robtobey any medals.